# Preprocessing Disks for Convex Hulls, Revisited

### v1

@@ 5.2.1 Fréchet and Hausdorff
The arrangement of all strips has compleixty $O(n)$
Proof 5.18.
@@
Proof 5.21 (Proof of Theorem 2.11 for disjoint disks).
@@
6 Preprocessing overlapping unit disks for convex hulls
We will now prove Theorems 2.10 and 2.11 in full generality for overlapping disks, by utilizing our one dimensional results, namely Lemma 3.7. Generally, the main ideas and structure of the arguments remain the same; hence we only describe the main differences.
Most importantly, Lemma 5.12 from Section 5 fails for overlapping disks, which implies that we need to construct our sequences differently. Instead we will reuse Lemma 3.7 from Section 3.2 as a subroutine.
@@ 6 Preprocessing overlapping unit disks for convex hulls
To prove Lemma 6.1 we will require a bound on the number of disks that can locally project onto a line.
@@
Proof 6.3.
We need to find a bound on the ply of the projected disks in $\cal I$. The maximum ply is given by $\Delta$ times the number of disjoint disks that fit inside a rectangle of dimensions $4\times(2d+2)$. Refer to Figure 10. Specht et al. [36] show that this number is at most $22\lceil d\rceil+$ 2.
Proof 6.4 (Proof of Lemma 6.1).
W.l.o.g. assume $\ell$ is horizontal. We will project all disks to $\ell$ and consider the resulting set of unit intervals $\cal I$. By Lemma 6.2, we can bound the ply of the projected disks in $\cal I$ by $\Delta^{\prime}=2\lceil d\rceil+2$. Then, we apply Lemma 3.7 to obtain a sequence $\Xi$ of at most $(8\Delta\lceil d\rceil+8\Delta)|\mathcal{R}|$ intervals, which translates directly to the required sequence of disks. Moreover, Lemma 3.7 also tells us that $\Xi$ is $(3,(4\Delta\lceil d\rceil+4\Delta))$-smooth.
@@
Consider the strip hull of a set $\mathcal{R}$ of unit disks of ply $\Delta$. Every unit disk intersects at most $714\Delta-$ 7 strips.
Proof 6.6.
Let $\cal B$ be the set of boundary disks of $\mathcal{R}$, and let $\cal I$ be a maximum independent set (in the intersection graph) of $\cal B$. By Corollary 5.3, every unit disk intersects at most $7$ strips of the strip hull of $\cal I$. Clearly every vertex of the strip hull of $\cal I$ is also a vertex of the strip hull of $\cal B$; furthermore, since every disk in $\cal B$ intersects at least one disk in $\cal I$, there can be at most $22\Delta-$ 2 additional vertices between any pair of adjacent vertices on the strip hull of $\cal B$; that is, each strip in of the strip hull of $\cal I$ corresponds to a sequence of at most $12\Delta-$ 1 strips in the strip hull of $\cal B$. The lemma follows.
@@
Consider the strip hull of a set $\mathcal{R}$ of unit disks of ply $\Delta$. The arrangement of strips has ply at most $48\Delta-$ 4.
Proof 6.8.
Let $\cal B$ and $\cal I$ be as in the proof of Lemma 6.5. By Corollary 5.5, Any point in the plane is contained in at most $4$ strips of the strip hull of $\cal I$, and thus in at most $48\Delta-$ 4 strips of the strip hull of $\cal B$.
Proof of Theorem 2.10.
@@
Furthermore, Lemma 6.1 states that when considering the intervals resulting from the projection of $\mathcal{R}$ on $\ell$, $\Xi_{i}$ is a $(3,(4\cdot 2+4)\Delta)=(3,12\Delta)$-smooth sorting-supersequence. Note that the angle between $\ell$ and a line of slope $-1$ is at most $45^{\circ}$; therefore, the distance $\alpha$ increases by at most a factor $\sqrt{2}$. Moreover, $\beta$ will not increase when going from the projection on the line back to the plane. We conclude that each $\Xi_{i}$ is $(3\sqrt{2},12\Delta)$-smooth.
Proof of Theorem 2.11.
@@
7 Reconstruction in Two Dimensions
In this section we present our reconstruction algorithm and then discuss how to extend it to achieve potential sublinear reconstruction by marking disks. Note the results in the section work directly for disks of bounded ply.
@@ 7.1 Standard Reconstruction
In this section, we consider the problem of recovering the true quarter hull from $\Xi$, given a realization $P\mathrel{\subset}\mkern-10.0mu\joinrel\sim\mathcal{R}$. We again start by replacing the disks in $\Xi$ by their corresponding points; now we have a sequence $X$ of points that contains the points on the quarter hull in counterclockwise order as a subsequence. Whether we can recover the quarter hull of such a sequence in linear time in general is equivalent to Open Problem 2.15. Here, we additionally assume the sequence $X$ is $(\alpha,\beta)$-smooth for constants $\alpha$ and $\beta$.
7.1.1 Algorithm for a Single Quadrant
@@
We now describe our algorithm for extracting the quarter hull. It is an adaptation of the algorithm of Andrew [5] (which is in turn an adaptation of Graham’s scan), but we don’t have a guarantee that the points come completely sorted. However, whenever we encounter a point that is not in the correct sorted order, we argue that we can determine whether it should still be inserted or can be safely discarded in $O(\alpha\cdot\beta)$ time be checking only the last $\alpha\cdot\beta$ edges of the hull so far, similar to the algorithm in Section 4.
@@
Proof 7.3.
We process the points in order, maintaining a convex chain, $C$, of a subset the points processed so far, where initially $C$ is just the rightmost point. So let $q$ be the next point to be processed, and let $p$ be the last (i.e. topmost) point on $C$. If $q=p$, then we skip $q$ and move on to processing the next point. Otherwise, If the $y$-coordinate of $q$ is greater than or equal to that of $p$, then we apply a standard Graham scan insert (i.e. pop from $C$ until it is a left turn to $q$, and then insert $q$ at the end of the chain). Finally, if the $y$-coordinate of $q$ is less than that of $p$ then there are two cases. Let $\ell$ be the line of slope $-1$ through $p$, and let $proj(q)$ be the orthogonal projection of $q$ onto $\ell$. If $||proj(q)-p||>\alpha$ then we pop $p$ from $C$, and we restart the process of trying to insert $q$. Otherwise, we check the last $\alpha\cdot\beta$ edges of $C$ in order from left to right, where if for any edge $q$ lies above the supporting line of that edge then we delete all vertices of $C$ starting from the left endpoint of that edge and going to the left, after which we then perform a Graham scan insert of $q$ into the remaining chain. If $q$ does not lie above any such edge, then we skip $q$ and move on to processing the next point.
@@
As for correctness, first observe that in the case that $||proj(q)-p||>\alpha$, then as $X$ is an $(\alpha,\beta)$-smooth sequence, by definition $p$ cannot be a vertex of the quarter hull. Thus the algorithm correctly deletes it, and so we will ignore this case in the remainder of the proof.
@@
Again let $C=c_{1},\ldots,c_{m}$ be the convex chain at some iteration of the algorithm. Suppose that the $i-1$ prefix of $C$, namely $C_{i-1}=c_{1},\ldots,c_{i-1}$, is the $i-1$ prefix of the quarter hull, and that the $i$th vertex of the quarter hull, call it $q$, is not in $c_{i},\ldots,c_{m}$. In this case, we claim that if $q$ is processed next by the algorithm then afterwards $C_{i}$ with be the $i$th prefix of the quarter hull. As prefixes of the quarter hull are never removed as argued above, proving this claim inductively implies we correctly compute the quarter hull, as we know the quarter hull is a subsequence of $X$ and thus we must eventually process the next vertex of the quarter hull. (The base case holds as initially $C$ is the first point in $X$, which by the lemma statement is the first vertex of the quarter hull.)
To prove the claim we consider the different cases handled by the algorithm, in order. First, $q=c_{m}$ ($c_{m}$ is called $p$ above) is not possible since $q$ by assumption of the claim statement is not in $C$. If the $y$-coordinate of $q$ is greater or equal to $c_{m}$ then we perform a Graham scan insert of $q$, which is guaranteed to place $q$ in $C$ after $c_{i-1}$ (since the fact that $q$ is above $c_{m}$ implies that $c_{1},\ldots,c_{i-1}$ are the only vertices of the quarter hull in $C$). So suppose that the $y$-coordinate of $q$ is less than that of $c_{m}$. Next, in the case that $||proj(q)-c_{m}||>\alpha$, we already remarked that $c_{m}$ is correctly deleted. So suppose that $||proj(q)-c_{m}||\leq\alpha$.
Now because $q$ is on the actual quarter hull, though it lies below $c_{m}$, it must lie to the right of $c_{m}$. In particular, it must lie above the chain $C$, i.e. in the region right of and below $c_{m}$, above and left of $c_{1}$, and above $C$. Thus $q$ lies above some edge of $C$, and since $||proj(q)-c_{m}||\leq\alpha$, as $X$ is an $(\alpha,\beta)$-smooth sequence, the packing property implies it must be one of the last $\alpha\cdot\beta$ edges of $C$. Finally, after identifying and removing up to the right end of this edge, $q$ will make a right turn with the remainder of $C$ until $c_{i-1}$. Thus after applying a Graham scan update, $q$ will be placed after $c_{i-1}$ as claimed.
@@
7.1.2 Combining the Pieces
Proof 7.4 (Proof of Theorem 2.12).
If $\Xi$ is an $(\alpha,\beta)$-smooth quarter-hull-supersequence of $\mathcal{R}$, then by definition for any $P\mathrel{\subset}\mkern-10.0mu\joinrel\sim\mathcal{R}$, if we replace the regions from $\Xi$ by their corresponding points in $P$, then it results in an $(\alpha,\beta)$-smooth sequence such that the quarter hull of $P$ is a subsequence. Moreover, by Observation 7.1.1 we can assume the first point in the sequence is the rightmost and the last is the topmost. Thus Lemma 7.2 directly states we can recover the quarter-hull of $P$ in $O(\alpha\beta|\Xi|)$ time.
@@ 7.2 Sublinear Reconstruction
An advantage of using supersequences as an auxiliary structures is that $\Xi^{\prime}$ for our problem is simply a sequence of points and disks, which is the desired output convex hull when replacing the disks in the sequence with their corresponding realizations. As this second subphase is now trivial, we focus on the first subphase in the remainder of this section. This requires identifying the disks whose realizations do not affect the combinatorial structure of the convex hull, which are precisely the stable disks defined in Section 2.1.
As argued in Section 2.1, there are two types of stable disks, namely stable guaranteed boundary disks and stable impossible interior disks. First observe that stable impossible interior disks do not properly intersect any strip on the boundary (that is $I$ from Observations 2.1 lies inside the chain of strips). Thus our constructed supersequence $\Xi$ already omits such disks, as it is constructed only from disks which intersect strips.
@@
We now adapt the preprocessing algorithm to build a marked supersequence $\Xi$ as follows. Let the counterclockwise sorted order of boundary disks be ()=$\mathcal{B}(\mathcal{R})=\{B_{1},\ldots,B_{k}\}$. The marked disks in this ordering intuitively partition the disks into maximal length substrings of unmarked disks, which in turn partitions the strip hull into subchains. In particular, if $B_{i}$ is a marked disk then both of its adjacent strips ($B_{i-1}B_{i}$ and $B_{i}B_{i+1}$) are no longer considered, which is the desired behavior as recall the proof of Lemma 2.6 argued no other disks intersected these strips, and because $B_{i}$ is stable the location of its realization does not affect the result on the substring ending at $B_{i-1}$ (and analogously for $B_{i+1}$). Thus we can safely apply Theorem 2.11 independently to each resulting substring and concatenate the results in order, with the respective marked disks in between.
@@ 8 Conclusion & Future Work
Figure 11: When we grow the uncertainty radii from 00 to $\infty$, all disks eventually turn from stable (red or green) to unstable potential (orange or yellow).
We introduced the idea of using supersequences as auxiliary structures in the preprocessing model for dealing with uncertain geometric data. We were able to reproduce several known results with an arguably simpler auxiliary structure; moreover, we obtained at least one new result.
@@
To follow up these first results using supersequences, we see three main lines of future work.
First, our work exposed an open problems that is of independent interest (Open Problem 2.15); settling whether this problem allows solutions in $o(n\log n)$ time or has a $\Omega(n\log n)$ lower bound would be very interesting.
Second, our results are restricted to unit intervals or disks of constant ply. Previous work in the preprocessing model has focused on allowing more general sets of regions, that may overlap more, be less fat, or have different sizes - which combinations of such generalizations are possible and for which problems is the topic of a whole line of papers. It would be very interesting to see whether our results can be generalized to less restricted regions as well; likely, such generalizations would require a relaxation of the smooth property, thus linking this line of future work to solving the open problems as well. Relatedly, when the uncertainty radius is not fixed in advance, understanding the relation between the radius, the resulting ply, and the efficiency of reconstruction would be of interest (see Figure 11).
@@
Finally, we only looked at two problems that produce an output that is a subsequence of the input: sorting and the convex hull. There are other natural problems that might fit in this framework, such as the Pareto Front, the traveling salesperson problem, shortest paths in polygons, and more. It would be interesting to study such problems in the same model.
Acknowledgements
The authors would like to thank Ivor van der Hoog, Vahideh Keikha, and Frank Staals for inspiring discussions on the topic of this paper.

### v2

@@ 4.2.2 Computation of a Supersequence for Disjoint Unit Disks
The arrangement of all strips has complexity $O(n)$
Proof 4.20.
@@
Proof 4.23 (Proof of Theorem 4.18).
@@
4.3 Preprocessing Overlapping Unit Disks
We will now prove Theorem 3.4 for overlapping unit disks.
Theorem 4.24.
Let $\mathcal{R}$ be a set of $n$ unit disks in the plane of ply $\Delta$. There exists a $(3\sqrt{2},12\Delta)$-smooth quarter-hull-supersequence $\Xi$ of $\mathcal{R}$ of size $O(\Delta^{2}n)$. $\Xi$ can be computed in $O(\Delta n(\Delta+\log n))$ time.
The main ideas and structure of the arguments remain the same; hence we only describe the main differences. Most importantly, Lemma 4.13 from Section 4.2 fails for overlapping disks, which implies that we need to construct our sequences differently. We will set out to prove the following.
@@ 4.3 Preprocessing Overlapping Unit Disks
In order to prove the above lemma, we need several components. First, we will need the following lemma concerning sorting supersequences in 1d from [32]:
Lemma 4.26 ([32][Lemma 13).
[…]
Second, to prove Lemma 4.25, we will also require a bound on the number of disks that can locally project onto a line.
@@
Proof 4.28.
[…]
Now, we are ready to prove Lemma 4.25.
Proof 4.29 (Proof of Lemma 4.25).
[…]
@@
Consider the strip hull of a set $\mathcal{R}$ of unit disks of ply $\Delta$. Every unit disk intersects at most $14\Delta-7$ strips.
Proof 4.31.
[…]
@@
Consider the strip hull of a set $\mathcal{R}$ of unit disks of ply $\Delta$. The arrangement of strips has ply at most $8\Delta-4$.
Proof 4.33.
[…]
Proof 4.34 (Proof of Theorem 4.24).
@@
Furthermore, Lemma 4.25 states that when considering the intervals resulting from the projection of $\mathcal{R}$ on $\ell$, $\Xi_{i}$ is a $(3,(4\cdot 2+4)\Delta)=(3,12\Delta)$-smooth sorting-supersequence. Note that the angle between $\ell$ and a line of slope 1-1 is at most $45^{\circ}$; therefore, the distance $\alpha$ increases by at most a factor $\sqrt{2}$. Moreover, $\beta$ will not increase when going from the projection on the line back to the plane. We conclude that each $\Xi_{i}$ is $(3\sqrt{2},12\Delta)$-smooth.
@@
4.4 Preprocessing Overlapping Disks of Varying Radii
Note that, unlike for overlapping disks, there is no reason a priori to expect that disks of arbitrary radii cannot be handled. However, the machinery developed is this paper relies heavily on the convexity of the spine hull, which we lose in this case (Figure 5 (c)).
Nonetheless, here we show that our results do extend to the case when the disks have varying radii when there is a bounded ratio $\rho$ between the smallest and largest disk.
The basic idea is that we can simply enlarge all disks about their centers so that they have a uniform radius of $\rho$. Doing so will potentially increase the ply, and we now upper bound this value. So consider any point $x$ in the plane. Observe that any disk that overlaps $x$ after enlargement must have its center within a ball of radius $\rho$ about $x$. As the enlargement does not move the disks centers, this implies any such disks must be entirely contained in a ball of radius $2\rho$ about $x$ before enlargement. A simple volumetric argument implies that the maximum number of unit balls of ply $\Delta$ that are entirely contained in a ball of radius $2\rho$ is at most $\Delta(\pi(2\rho)^{2}/\pi(1)^{2})=4\Delta\rho^{2}$. (A more careful argument should improve the factor of 4.)
Theorem 4.35.
Let $\mathcal{R}$ be a set of $n$ disks in the plane, with minimum radius $1$ and maximum radius $\rho$, and let $\Delta$ be the ply of $\mathcal{R}$. A $(O(1),O(\rho^{2}\Delta))$-smooth quarter-hull-supersequence $\Xi$ of $\mathcal{R}$ of size $O(\rho^{4}\Delta^{2}n)$ can be computed in $O(\rho^{2}\Delta n(\rho^{2}\Delta+\log n))$ time.
5 The Reconstruction Phase
In this section we present our reconstruction algorithm and then discuss how to extend it to achieve potential sublinear reconstruction by marking disks. Note the results in this section do not require the disks be unit disks nor that they have bounded ply. Indeed, we only know how to construct an $(\alpha,\beta)$-smooth supersequence for unit disks of bounded ply; however, we believe it is still of value to have a reconstruction algorithm for the more general case. We will first consider standard reconstruction (that is, aiming for linear time) in Section 5.1, and then discuss sublinear reconstruction in Section 5.2.
@@ 5.1 Standard Reconstruction
In this section, we consider the problem of recovering the true quarter hull from $\Xi$, given a realization $P\mathrel{\subset}\mkern-10.0mu\joinrel\sim\mathcal{R}$. We again start by replacing the disks in $\Xi$ by their corresponding points; now we have a sequence $X$ of points that contains the points on the quarter hull in counterclockwise order as a subsequence. Whether we can recover the quarter hull of such a sequence in linear time in general is equivalent to Open Problem 1.2. Here, we additionally assume the sequence $X$ is $(\alpha,\beta)$-smooth for constants $\alpha$ and $\beta$.
Algorithm for a Single Quadrant
@@
We now describe our algorithm for extracting the quarter hull. It is an adaptation of the algorithm of Andrew [3] (which is in turn an adaptation of Graham’s scan), but we don’t have a guarantee that the points come completely sorted. However, whenever we encounter a point that is not in the correct sorted order, we argue that we can determine whether it should still be inserted or can be safely discarded in $O(\alpha\cdot\beta)$ time be checking only the last $\alpha\cdot\beta$ edges of the hull so far.
@@
Proof 5.3.
We process the points in order, maintaining a convex chain, $C$, of a subset the points processed so far, where initially $C$ is just the rightmost point. So let $q$ be the next point to be processed, and let $p$ be the last (i.e. topmost) point on $C$. If $q=p$, then we skip $q$ and move on to processing the next point. Otherwise, If the $y$-coordinate of $q$ is greater than or equal to that of $p$, then we apply a standard Graham scan insert (i.e. pop from $C$ until it is a left turn to $q$, and then insert $q$ at the end of the chain). Finally, if the $y$-coordinate of $q$ is less than that of $p$ then there are two cases. Let $\ell$ be the line of slope 1-1 through $p$, and let $proj(q)$ be the orthogonal projection of $q$ onto $\ell$. If $proj(q)-p||>\alpha$ then we pop $p$ from $C$, and we restart the process of trying to insert $q$. Otherwise, we check the last $\alpha\cdot\beta$ edges of $C$ in order from left to right, where if for any edge $q$ lies above the supporting line of that edge then we delete all vertices of $C$ starting from the left endpoint of that edge and going to the left, after which we then perform a Graham scan insert of $q$ into the remaining chain. If $q$ does not lie above any such edge, then we skip $q$ and move on to processing the next point.
@@
As for correctness, first observe that in the case that $proj(q)-p||>\alpha$, then as $X$ is an $(\alpha,\beta)$-smooth sequence, by definition $p$ cannot be a vertex of the quarter hull. Thus the algorithm correctly deletes it, and so we will ignore this case in the remainder of the proof.
@@
Again let $C=c_{1},\ldots,c_{m}$ be the convex chain at some iteration of the algorithm. Suppose that the i1i-1 prefix of $C$, namely $C_{i-1}=c_{1},\ldots,c_{i-1}$, is the i1i-1 prefix of the quarter hull, and that the $i$th vertex of the quarter hull, call it $q$, is not in $c_{i},\ldots,c_{m}$. In this case, we claim that if $q$ is processed next by the algorithm then afterwards $C_{i}$ with be the $i$th prefix of the quarter hull. As prefixes of the quarter hull are never removed as argued above, proving this claim inductively implies we correctly compute the quarter hull, as we know the quarter hull is a subsequence of $X$ and thus we must eventually process the next vertex of the quarter hull. (The base case holds as initially $C$ is the first point in $X$, which by the lemma statement is the first vertex of the quarter hull.)
To prove the claim we consider the different cases handled by the algorithm, in order. First, $q=c_{m}$ ($c_{m}$ is called $p$ above) is not possible since $q$ by assumption of the claim statement is not in $C$. If the $y$-coordinate of $q$ is greater or equal to $c_{m}$ then we perform a Graham scan insert of $q$, which is guaranteed to place $q$ in $C$ after $c_{i-1}$ (since the fact that $q$ is above $c_{m}$ implies that $c_{1},\ldots,c_{i-1}$ are the only vertices of the quarter hull in $C$). So suppose that the $y$-coordinate of $q$ is less than that of $c_{m}$. Next, in the case that $proj(q)-c_{m}||>\alpha$, we already remarked that $c_{m}$ is correctly deleted. So suppose that $proj(q)-c_{m}||\leq\alpha$.
Now because $q$ is on the actual quarter hull, though it lies below $c_{m}$, it must lie to the right of $c_{m}$. In particular, it must lie above the chain $C$, i.e. in the region right of and below $c_{m}$, above and left of $c_{1}$, and above $C$. Thus $q$ lies above some edge of $C$, and since $proj(q)-c_{m}||\leq\alpha$, as $X$ is an $(\alpha,\beta)$-smooth sequence, the packing property implies it must be one of the last $\alpha\cdot\beta$ edges of $C$. Finally, after identifying and removing up to the right end of this edge, $q$ will make a right turn with the remainder of $C$ until $c_{i-1}$. Thus after applying a Graham scan update, $q$ will be placed after $c_{i-1}$ as claimed.
@@
Combining the Pieces
Proof 5.4 (Proof of Theorem 3.5).
If $\Xi$ is an $(\alpha,\beta)$-smooth quarter-hull-supersequence of $\mathcal{R}$, then by definition for any $P\mathrel{\subset}\mkern-10.0mu\joinrel\sim\mathcal{R}$, if we replace the regions from $\Xi$ by their corresponding points in $P$, then it results in an $(\alpha,\beta)$-smooth sequence such that the quarter hull of $P$ is a subsequence. Moreover, by Observation 5.1 we can assume the first point in the sequence is the rightmost and the last is the topmost. Thus Lemma 5.2 directly states we can recover the quarter-hull of $P$ in $O(\alpha\beta|\Xi|)$ time.
@@ 5.2 Sublinear Reconstruction
An advantage of using supersequences as an auxiliary structures is that $\Xi^{\prime}$ for our problem is simply a sequence of points and disks, which is the desired output convex hull when replacing the disks in the sequence with their corresponding realizations. As this second subphase is now trivial, we focus on the first subphase in the remainder of this section. This requires identifying the disks whose realizations do not affect the combinatorial structure of the convex hull, which are precisely the stable disks defined in Section 2.
As argued in Section 2, there are two types of stable disks, namely stable guaranteed boundary disks and stable impossible interior disks. First observe that stable impossible interior disks do not properly intersect any strip on the boundary (that is $I$ from Observations 2 lies inside the chain of strips). Thus our constructed supersequence $\Xi$ already omits such disks, as it is constructed only from disks which intersect strips.
@@
We now adapt the preprocessing algorithm to build a marked supersequence $\Xi$ as follows. Let the counterclockwise sorted order of boundary disks be $\mathcal{B}(\mathcal{R})=\{B_{1},\ldots,B_{k}\}$. The marked disks in this ordering intuitively partition the disks into maximal length substrings of unmarked disks, which in turn partitions the strip hull into subchains. In particular, if $B_{i}$ is a marked disk then both of its adjacent strips ($B_{i-1}B_{i}$ and $B_{i}B_{i+1}$) are no longer considered, which is the desired behavior as recall the proof of Lemma 2.8 argued no other disks intersected these strips, and because $B_{i}$ is stable the location of its realization does not affect the result on the substring ending at $B_{i-1}$ (and analogously for $B_{i+1}$). Thus we can safely apply Theorem 3.4 independently to each resulting substring and concatenate the results in order, with the respective marked disks in between.
@@ 6 Conclusion & Future Work
We introduced the idea of using supersequences as auxiliary structures in the preprocessing model for dealing with uncertain geometric data. To follow up these first results using supersequences, we see three main lines of future work.
First, our preprocessing results are currently restricted to bounded radius ratio disks of constant ply. In principle, it should be possible to generalize these results, e.g. to arbitrary disks, or fat regions, although most likely the smooth property in its current form would be too restrictive.
Second, our work exposed Open Problem 1.2, which we believe is of independent interest, but also has direct implications on the preprocessing framework: a linear-time solution to this problem would immediately resolve the reconstruction problem on any supersequence, independent of the type of regions or any smoothness condition.
@@
Finally, we only looked at constructing a supersequence for the convex hull. There are other natural problems that might fit in this framework, such as the Pareto Front, the traveling salesperson problem, shortest paths in polygons, and more. It would be interesting to study such problems in the same model.
Appendix A Prior Work on Uncertain Convex Hulls
The convex hull is one of the oldest and most studied structures in computational geometry, dating back to Graham [19]; an overview of the literature on convex hull algorithms is beyond the scope of this work. Within the relatively younger field of uncertain computational geometry, the convex hull is still one of the most-studied problems.
Boissonnat and Lazard [4] study the problem of finding the shortest convex hull of bounded curvature that contains a set of points, and they show that this is equivalent to finding the shortest convex hull of a set of imprecise points modelled as circles that have the specified curvature. They give a polynomial-time approximation algorithm. Goodrich and Snoeyink [18] also show how to minimize the area or perimeter of the polygon in $O(n^{2})$ time.
Van Kreveld and Löffler [30] study the problem of computing upper and lower bounds on the area or perimeter of the convex hull. Ju and Luo [26] improve one of these results, and also consider some variations in the model of imprecision. Some of the resulting problems here have also been studied in a different context. Mukhopadhyay et al. [34, 33] study the largest area convex polygonal stabber of a set of parallel or isothetic line segments, and similarly, Hassanzadeh and Rappaport [21, 39] study the shortest perimeter convex polygonal stabber of a set of line segments.
Many geometric algorithms return a certain region in the plane, or a subdivision of the plane into regions. In these cases, a natural way of representing imprecise output is by providing two boundaries of such a region: an inner boundary that encloses all points that are certain to be in the region, and an outer boundary that encloses all points that could be in the region. Such a pair of boundaries is sometimes referred to as a fuzzy boundary [41]. Nagai and Tokura [35] follow this approach for the convex hull by computing the union and intersection of all possible convex hulls. As imprecision regions they use disks and convex polygons, and they give an $O(n\log n)$ time algorithm for computing both boundaries. Sember and Evans [15, 42] define similar concepts.
Bruce et al. [5] consider update strategies for convex hulls for moving points, which is very similar to the static case with uncertainty. They introduce (inspired in turn by Feder et al. [17]) a notion of uncertain points being always, never, or dependent on the convex hull, which is similar to part of our classification in Section 2. However, their focus is on the number of retrievals rather than computation time: they assume an oracle is available which can be used to retrieve, for any region, the true location of its point, and aim to develop a strategy which, for any given instance, performs the smallest number of retrievals possible and still guarantees the correct output. Van der Hoog et al. [47] later term such a strategy to be instance optimal if it performs no more retrievals than necessary by comparing to an adversarial strategy, though they focus on the Pareto front and not the convex hull. Very recently, de Berg et al. [9] show how to do this in instance optimal time for the convex hull as well.